# Towards Black-Box Membership Inference Attack for Diffusion Models

### v1

@@ Appendix C Proof
If we denote the random variable $X_{i}$ represents $\|\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t))\|$, then we consider $S_{n}=\sum_{i=1}^{n}X_{i}$. From the assumption we know that all the $X_{i}$ are centered random variable $X$ with finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$. Using the theorem of Chernoff-Cramer method for sums of IID RV [7], we get the following probability inequality for any $\beta>0$:
$\mathbb{P}(S_{n}\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta}{n}))\,.$
Therefore, taking the definition of $\|\hat{x}-x\|=\|\frac{1}{n}\sum_{i=1}^{n}(x^{i}-x)\|\leq\frac{\sqrt{1-\bar{\alpha}_{t}}}{n\sqrt{\bar{\alpha}_{t}}}S_{n}$, we get the following bound of the reconstruction error:
$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}))\,,$
where $\beta>0$ and $\Psi^{*}_{X}(y)=\sup_{s>0}(sy-\Psi_{X}(s))$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$. So averaging the randomly reconstructed data from a training set will result in a smaller reconstruction error with high probability of $\Theta(1-\exp(-n))$ when we use a large $n$.

### v2

@@ Appendix C Proof
If we denote the random variable $X^{i}=(X^{i}_{1},X^{i}_{2},\ldots,X^{i}_{d})$ represents $\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)$, then we consider $S^{i}_{n}=\sum_{j=1}^{n}X^{i}_{j}$. From the assumption we know that all the $X^{i}_{j}$ are random variables with zero expectation and finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$. Using the theorem of Chernoff-Cramer method for sums of IID RV [8], we get the following probability inequality for any $\beta>0$:
$\mathbb{P}(|S^{i}_{n}|\geq\beta)\leq\exp(-n\Psi^{*}_{X_{i}}(\frac{\beta}{n}))\,,$
where $\beta>0$ and $\Psi^{*}_{X}(y)=\sup_{s>0}(sy-\Psi_{X}(s))$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$.
Therefore, denote $S_{n}=(S^{1}_{n},S^{2}_{n},\ldots,S^{d}_{n})$, taking the definition of $\|\hat{x}-x\|=\|\frac{1}{n}\sum_{i=1}^{n}(x^{i}-x)\|=\frac{\sqrt{1-\bar{\alpha}_{t}}}{n\sqrt{\bar{\alpha}_{t}}}\|S_{n}\|$, we get the following bound of the reconstruction error:
$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\sum^{d}_{i=1}P(|S^{i}_{n}|\geq\frac{n\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{d(1-\bar{\alpha}_{t}})})\leq d\exp(-n\min_{i}\Psi^{*}_{X_{i}}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{d(1-\bar{\alpha}_{t}})}))\,,$
So averaging the randomly reconstructed data from a training set will result in a smaller reconstruction error with high probability of $\Theta(1-\exp(-n))$ when we use a large $n$.
Appendix D More Results of Variation Images
In this section, we provide more results of the variation of member and nonmember image when applying to DALL-E 2’s variation API. The experimental results are recorded in Figure 16, from which we can see that the variation in the member inputs after applying the DALL-E 2’s variation API is relatively small than non-member inputs.
Figure 16: More results of the variation images. From the figures we can see that DALL-E 2’s variation API makes less changes to images in member set than non-member set.